When 50.L of neon gas at 25 celsius compressed to 25.L, what is the new temperature of neon gas?
Answer #1

Given:

Ti = 25.0 oC

= (25.0+273) K

= 298 K

Vi = 50 L

Vf = 25 L

use:

Vi/Ti = Vf/Tf

50.0 L / 298.0 K = 25.0 L / Tf

Tf = 149 K

Answer:  149 K